# Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling

### v2

@@ 3.5 Final Estimation Setup
Following Cont and Das, (2024), we apply the estimator directly to the series of realised volatility magnitudes, defined as $RV_{t}=|r_{t}|$, where $r_{t}$ is the log-return at time $t$. This approach avoids any transformation to log-volatility or its increments, and the resulting roughness index $\hat{H}$ captures the fine-scale regularity of the volatility path itself. Summary statistics of $\log W(L,K,p)$ behaviour are reported in Table 3 in Section 4.1.
@@ 3.6.1 Non-Stationarity Diagnostics
To assess whether the failure of the $p$-variation estimator may stem from non-stationarity or structural breaks, we apply three complementary diagnostics to each realised volatility series: the Augmented Dickey-Fuller (ADF) test for mean stationarity, rolling window analysis to evaluate local stability of mean and variance, and structural break detection to identify discrete regime shifts. Together, these methods provide a comprehensive assessment of both gradual and abrupt departures from stationarity.
@@ 3.6.2 Multifractality Diagnostics
To assess whether Bitcoin’s realised volatility exhibits a multifractal temporal structure, we apply three complementary diagnostics: Multifractal Detrended Fluctuation Analysis (MF-DFA), log-log moment scaling, and wavelet leaders analysis. Each method is implemented systematically across all (year, frequency) combinations and applied uniformly to both log-return and realised volatility series. This convention is maintained throughout and not repeated in the descriptions that follow.
@@
is computed, where $F_{q}(s)$ denotes the $q$th-order fluctuation function at scale $s$. Summary statistics for the resulting $H(q)$ spectra (including minimum, maximum, mean, and spectral width) are reported in Appendix Table A.6 (log-returns) and Table A.9 (realised volatility).
@@
is computed, where $M_{q}(\tau)$ denotes the $q$th-order moment of increments at lag $\tau$. This procedure is implemented over a dense grid of lags $\tau$. Complete results are presented in Appendix Table A.7 (log-returns) and Table A.10 (realised volatility).
@@
is estimated, where $M_{q}(2^{j})$ denotes the $q$th-order moment of wavelet leaders at dyadic scale $2^{j}$. Scaling exponents $\zeta_{q}$ are obtained via linear regression, and corresponding $R^{2}$ values are recorded to evaluate the quality of fit. Wavelet leaders, constructed from the local suprema of wavelet coefficients across scales, offer robustness to non-stationarities and facilitate the isolation of multiscale fluctuations from deterministic trends. Full results are presented in Appendix Table A.8 (log-returns) and Table A.11 (realised volatility).
@@ 4.1 Main Result: Roughness Estimation behaviour
Figure 1: Estimated roughness index $\widehat{H}^{L,K}$ versus number of blocks $K$ for BTC realised volatility in 2021 at 1-minute (top-left), 5-minute (top-right), 10-minute (bottom-left), and 15-minute (bottom-right) frequencies.
@@ 4.2 Stationarity and Window Stability
Diagnostic results do not support non-stationarity or structural breaks as plausible explanations for the failure of the normalised $p$-variation estimator. Bitcoin realised volatility appears stationary with stable local dynamics. If non-stationarity were the cause, one would expect to observe systematic trends, level shifts, or persistent regime changes, none of which are observed in the diagnostic results.
@@ 4.3 Evidence of Multifractality
The results of the Multifractal Detrended Fluctuation Analysis (MF-DFA), described in Section 3.6.2, are reported in Appendix Tables A.6 and A.9. Across all original series, the estimated $H(q)$ spectra exhibit pronounced non-linearity with respect to $q$, indicating scale-dependent behaviour characteristic of multifractality. For realised volatility, $H(q)$ remain consistently below 1 across all years and sampling frequencies, typically centred around 0.9 for moderate $q$, with spectral widths between 0.06 and 0.42. Log-returns display lower mean exponents but retain comparable curvature in the scaling function. When applied to shuffled series, removing temporal structure while preserving the marginal distribution, the $H(q)$ spectra compress substantially, with spectral widths falling below 0.15 in most cases. This contrast supports the interpretation that the observed multifractality arises from temporal dependence rather than distributional effects.
@@
We next apply the log-log moment scaling analysis, an alternative multifractality diagnostic based on the scaling behaviour of $\zeta_{q}$, as described in Section 3.6.2. Estimated $\zeta_{q}$ spectra for realised volatility and log-returns are presented in Appendix Tables A.7 and A.10. For realised volatility, the $\zeta_{q}$ spectra exhibit pronounced curvature in $q$, with widths ranging from 0.04 to over 2.0 and mean $R^{2}$ values generally exceeding 0.85, indicating stable scaling across temporal resolutions. Log-return series show more heterogeneity (especially at lower sampling frequencies) but maintain clear non-linearity in the $\zeta_{q}$ profile for moderate to large $q$. In the shuffled series, the $\zeta_{q}$ spectra flatten substantially, with widths typically below 0.2 and minimal curvature. These patterns are consistent with a temporal origin of the observed multifractality in both series.
@@
The wavelet leaders multifractal analysis offers a robust characterisation of scaling behaviour, particularly in the presence of trends or non-stationarities. Estimated $\zeta_{q}$ exponents for both realised volatility and log-return series are presented in Appendix Tables A.8 and A.11. For realised volatility, the $\zeta_{q}$ spectra exhibit consistent curvature with spectral widths between 0.04 and 0.51 and mean $R^{2}$ values typically exceeding 0.85. Log-return series show similar but generally weaker curvature, with greater sensitivity to sampling resolution. In the shuffled series, spectral widths decline substantially (often by more than 50%) and curvature is notably reduced.
@@
The MF-DFA, log-log moment scaling, and wavelet leaders analyses consistently indicate that Bitcoin realised volatility exhibits a multifractal temporal structure. This conclusion is supported by the marked reduction in spectral curvature and width observed in shuffled controls, which preserve the marginal distribution but remove temporal dependence. These results suggest that the observed multifractality arises from temporal correlations rather than from heavy-tailed marginal distributions. This violates the homogeneity assumptions required by the normalised $p$-variation framework and provides a structural explanation for the failure to identify a well-defined scaling exponent. The apparent roughness at small $q$ reflects localised scaling properties embedded in a broader multifractal structure, rather than a stable global Hurst exponent.
@@ 4.4 Hypothesis Evaluation
Our primary empirical result, the strict negativity of $\log W(L,K,p)$ across all years, frequencies, and $p$-grids, demonstrates the absence of any stable scaling law compatible with the normalised $p$-variation estimator in Bitcoin realised volatility. This supports the interpretation that the apparent roughness reported in prior literature does not reflect an intrinsic property of the volatility process, but rather the failure of the $p$-variation framework to capture consistent scaling behaviour in this context. These findings confirm Hypothesis H1, as the strict negativity of $\log W(L,K,p)$ across all configurations indicates the absence of a well-defined scaling law.
@@
Our suite of stationarity diagnostics, including the Augmented Dickey-Fuller test for unit roots, rolling mean and variance statistics for local stability, and structural break detection via binary segmentation with an $L_{2}$-based cost function, yields no evidence that non-stationarity or persistent structural instability accounts for the failure of the $p$-variation estimator. The realised volatility series are stationary over the studied periods, and the detected breakpoints correspond to local volatility clustering rather than structural regime changes. We therefore reject $H_{3}$ as a primary explanation for the observed behaviour of $\log W(L,K,p)$, given the absence of non-stationarity or structural instability.
@@
The multifractality hypothesis is decisively supported by our results. All three diagnostic approaches, MF-DFA, log-log moment scaling, and wavelet leaders analysis, exhibit consistent non-linear scaling in the original series and show substantial collapse of multifractal indicators under shuffle-based controls. This confirms that the observed multifractality in Bitcoin volatility, particularly in realised volatility, arises from genuine temporal correlations rather than from heavy-tailed marginal distributions. The presence of this multifractal temporal structure implies non-linear, scale-dependent behaviour that directly violates the homogeneity assumption of the normalised $p$-variation framework. We therefore accept $H_{4}$ as the most plausible explanation for the observed behaviour of $\log W(L,K,p)$, as the multifractal structure of Bitcoin volatility is supported by all diagnostic results and offers a coherent account of the estimator’s failure.
@@ 5.1 Structural Interpretation
Our analysis demonstrates that Bitcoin realised volatility exhibits pronounced scale dependence and temporal heterogeneity, in contrast to the patterns typically observed in traditional financial assets. Evidence from multiple diagnostics, reinforced by shuffled controls, consistently indicates the presence of multifractality and the absence of a global scaling regime. This empirical structure is incompatible with the assumption of homogeneous Hölder continuity underlying the rough volatility framework.
@@ 5.4 Limitations and Future Work
While this study highlights the limitations of rough volatility estimation for Bitcoin and confirms the presence of multifractal structure, several limitations should be acknowledged. The study focuses exclusively on realised volatility derived from high-frequency price data. Although well-suited for $p$-variation analysis, this measure captures only a subset of the latent volatility process. The scaling behaviour of alternative measures, such as implied volatility or state-space estimates, has not been examined. Extending the analysis to these measures would clarify whether multifractality is an intrinsic property of Bitcoin volatility or an artefact of realised volatility estimation, and would offer insights into how market expectations reflect multiscale behaviour.
@@ 6 Conclusion
This study assessed the applicability of rough volatility methodologies to Bitcoin using the normalised $p$-variation framework proposed by Cont and Das, (2024). The statistic $\log W(L,K,p)$ remains strictly negative across all configurations, precluding the estimation of a valid roughness index. Diagnostic tests rule out non-stationarity, structural breaks, and sampling artefacts as plausible explanations. Instead, consistent evidence from MF-DFA, log-log moment scaling, and wavelet leaders indicates that Bitcoin realised volatility follows a robust multifractal temporal structure, driven by dynamic correlations rather than marginal distributional effects. This violates the homogeneity assumptions required by rough volatility models and accounts for the systematic failure of the estimator.
@@
Table A.9: Summary of MF-DFA Scaling Exponents $H(q)$ for Bitcoin Realised Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of $H(q)$ spectra.
@@
Table A.10: Summary of Log-Log Scaling Exponents for Bitcoin Realised Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of scaling coefficients.
@@
Table A.11: Summary of Wavelet Leaders Scaling Exponents for Bitcoin Realised Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) for each case.
@@
Table A.12: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, using the noise-robust realised volatility $\widetilde{RV}^{\Delta}_{t}$.

### v3

@@ 3.5 Final Estimation Setup
Following Cont and Das, (2024), we apply the estimator directly to the series of realized volatility magnitudes, defined as $RV_{t}=|r_{t}|$, where $r_{t}$ is the log-return at time $t$. This approach avoids any transformation to log-volatility or its increments, and the resulting roughness index $\hat{H}$ captures the fine-scale regularity of the volatility path itself. Summary statistics of $\log W(L,K,p)$ behaviour are reported in Table 3 in Section 4.1.
@@ 3.6.1 Non-Stationarity Diagnostics
To assess whether the failure of the $p$-variation estimator may stem from non-stationarity or structural breaks, we apply three complementary diagnostics to each realized volatility series: the Augmented Dickey-Fuller (ADF) test for mean stationarity, rolling window analysis to evaluate local stability of mean and variance, and structural break detection to identify discrete regime shifts. Together, these methods provide a comprehensive assessment of both gradual and abrupt departures from stationarity.
@@ 3.6.2 Multifractality Diagnostics
To assess whether Bitcoin’s realized volatility exhibits a multifractal temporal structure, we apply three complementary diagnostics: Multifractal Detrended Fluctuation Analysis (MF-DFA), log-log moment scaling, and wavelet leaders analysis. Each method is implemented systematically across all (year, frequency) combinations and applied uniformly to both log-return and realized volatility series. This convention is maintained throughout and not repeated in the descriptions that follow.
@@
is computed, where $F_{q}(s)$ denotes the $q$th-order fluctuation function at scale $s$. Summary statistics for the resulting $H(q)$ spectra (including minimum, maximum, mean, and spectral width) are reported in Appendix Table A.6 (log-returns) and Table A.9 (realized volatility).
@@
is computed, where $M_{q}(\tau)$ denotes the $q$th-order moment of increments at lag $\tau$. This procedure is implemented over a dense grid of lags $\tau$. Complete results are presented in Appendix Table A.7 (log-returns) and Table A.10 (realized volatility).
@@
is estimated, where $M_{q}(2^{j})$ denotes the $q$th-order moment of wavelet leaders at dyadic scale $2^{j}$. Scaling exponents $\zeta_{q}$ are obtained via linear regression, and corresponding $R^{2}$ values are recorded to evaluate the quality of fit. Wavelet leaders, constructed from the local suprema of wavelet coefficients across scales, offer robustness to non-stationarities and facilitate the isolation of multiscale fluctuations from deterministic trends. Full results are presented in Appendix Table A.8 (log-returns) and Table A.11 (realized volatility).
@@ 4.1 Main Result: Roughness Estimation behaviour
Figure 1: Estimated roughness index $\widehat{H}^{L,K}$ versus number of blocks $K$ for BTC realized volatility in 2021 at 1-minute (top-left), 5-minute (top-right), 10-minute (bottom-left), and 15-minute (bottom-right) frequencies.
@@ 4.2 Stationarity and Window Stability
Diagnostic results do not support non-stationarity or structural breaks as plausible explanations for the failure of the normalised $p$-variation estimator. Bitcoin realized volatility appears stationary with stable local dynamics. If non-stationarity were the cause, one would expect to observe systematic trends, level shifts, or persistent regime changes, none of which are observed in the diagnostic results.
@@ 4.3 Evidence of Multifractality
The results of the Multifractal Detrended Fluctuation Analysis (MF-DFA), described in Section 3.6.2, are reported in Appendix Tables A.6 and A.9. Across all original series, the estimated $H(q)$ spectra exhibit pronounced non-linearity with respect to $q$, indicating scale-dependent behaviour characteristic of multifractality. For realized volatility, $H(q)$ remain consistently below 1 across all years and sampling frequencies, typically centred around 0.9 for moderate $q$, with spectral widths between 0.06 and 0.42. Log-returns display lower mean exponents but retain comparable curvature in the scaling function. When applied to shuffled series, removing temporal structure while preserving the marginal distribution, the $H(q)$ spectra compress substantially, with spectral widths falling below 0.15 in most cases. This contrast supports the interpretation that the observed multifractality arises from temporal dependence rather than distributional effects.
@@
We next apply the log-log moment scaling analysis, an alternative multifractality diagnostic based on the scaling behaviour of $\zeta_{q}$, as described in Section 3.6.2. Estimated $\zeta_{q}$ spectra for realized volatility and log-returns are presented in Appendix Tables A.7 and A.10. For realized volatility, the $\zeta_{q}$ spectra exhibit pronounced curvature in $q$, with widths ranging from 0.04 to over 2.0 and mean $R^{2}$ values generally exceeding 0.85, indicating stable scaling across temporal resolutions. Log-return series show more heterogeneity (especially at lower sampling frequencies) but maintain clear non-linearity in the $\zeta_{q}$ profile for moderate to large $q$. In the shuffled series, the $\zeta_{q}$ spectra flatten substantially, with widths typically below 0.2 and minimal curvature. These patterns are consistent with a temporal origin of the observed multifractality in both series.
@@
The wavelet leaders multifractal analysis offers a robust characterisation of scaling behaviour, particularly in the presence of trends or non-stationarities. Estimated $\zeta_{q}$ exponents for both realized volatility and log-return series are presented in Appendix Tables A.8 and A.11. For realized volatility, the $\zeta_{q}$ spectra exhibit consistent curvature with spectral widths between 0.04 and 0.51 and mean $R^{2}$ values typically exceeding 0.85. Log-return series show similar but generally weaker curvature, with greater sensitivity to sampling resolution. In the shuffled series, spectral widths decline substantially (often by more than 50%) and curvature is notably reduced.
@@
The MF-DFA, log-log moment scaling, and wavelet leaders analyses consistently indicate that Bitcoin realized volatility exhibits a multifractal temporal structure. This conclusion is supported by the marked reduction in spectral curvature and width observed in shuffled controls, which preserve the marginal distribution but remove temporal dependence. These results suggest that the observed multifractality arises from temporal correlations rather than from heavy-tailed marginal distributions. This violates the homogeneity assumptions required by the normalised $p$-variation framework and provides a structural explanation for the failure to identify a well-defined scaling exponent. The apparent roughness at small $q$ reflects localised scaling properties embedded in a broader multifractal structure, rather than a stable global Hurst exponent.
@@ 4.4 Hypothesis Evaluation
Our primary empirical result, the strict negativity of $\log W(L,K,p)$ across all years, frequencies, and $p$-grids, demonstrates the absence of any stable scaling law compatible with the normalised $p$-variation estimator in Bitcoin realized volatility. This supports the interpretation that the apparent roughness reported in prior literature does not reflect an intrinsic property of the volatility process, but rather the failure of the $p$-variation framework to capture consistent scaling behaviour in this context. These findings confirm Hypothesis H1, as the strict negativity of $\log W(L,K,p)$ across all configurations indicates the absence of a well-defined scaling law.
@@
Our suite of stationarity diagnostics, including the Augmented Dickey-Fuller test for unit roots, rolling mean and variance statistics for local stability, and structural break detection via binary segmentation with an $L_{2}$-based cost function, yields no evidence that non-stationarity or persistent structural instability accounts for the failure of the $p$-variation estimator. The realized volatility series are stationary over the studied periods, and the detected breakpoints correspond to local volatility clustering rather than structural regime changes. We therefore reject $H_{3}$ as a primary explanation for the observed behaviour of $\log W(L,K,p)$, given the absence of non-stationarity or structural instability.
@@
The multifractality hypothesis is decisively supported by our results. All three diagnostic approaches, MF-DFA, log-log moment scaling, and wavelet leaders analysis, exhibit consistent non-linear scaling in the original series and show substantial collapse of multifractal indicators under shuffle-based controls. This confirms that the observed multifractality in Bitcoin volatility, particularly in realized volatility, arises from genuine temporal correlations rather than from heavy-tailed marginal distributions. The presence of this multifractal temporal structure implies non-linear, scale-dependent behaviour that directly violates the homogeneity assumption of the normalised $p$-variation framework. We therefore accept $H_{4}$ as the most plausible explanation for the observed behaviour of $\log W(L,K,p)$, as the multifractal structure of Bitcoin volatility is supported by all diagnostic results and offers a coherent account of the estimator’s failure.
@@ 5.1 Structural Interpretation
Our analysis demonstrates that Bitcoin realized volatility exhibits pronounced scale dependence and temporal heterogeneity, in contrast to the patterns typically observed in traditional financial assets. Evidence from multiple diagnostics, reinforced by shuffled controls, consistently indicates the presence of multifractality and the absence of a global scaling regime. This empirical structure is incompatible with the assumption of homogeneous Hölder continuity underlying the rough volatility framework.
@@ 5.4 Limitations and Future Work
While this study highlights the limitations of rough volatility estimation for Bitcoin and confirms the presence of multifractal structure, several limitations should be acknowledged. The study focuses exclusively on realized volatility derived from high-frequency price data. Although well-suited for $p$-variation analysis, this measure captures only a subset of the latent volatility process. The scaling behaviour of alternative measures, such as implied volatility or state-space estimates, has not been examined. Extending the analysis to these measures would clarify whether multifractality is an intrinsic property of Bitcoin volatility or an artefact of realized volatility estimation, and would offer insights into how market expectations reflect multiscale behaviour.
@@ 6 Conclusion
This study assessed the applicability of rough volatility methodologies to Bitcoin using the normalised $p$-variation framework proposed by Cont and Das, (2024). The statistic $\log W(L,K,p)$ remains strictly negative across all configurations, precluding the estimation of a valid roughness index. Diagnostic tests rule out non-stationarity, structural breaks, and sampling artefacts as plausible explanations. Instead, consistent evidence from MF-DFA, log-log moment scaling, and wavelet leaders indicates that Bitcoin realized volatility follows a robust multifractal temporal structure, driven by dynamic correlations rather than marginal distributional effects. This violates the homogeneity assumptions required by rough volatility models and accounts for the systematic failure of the estimator.
@@
Table A.9: Summary of MF-DFA Scaling Exponents $H(q)$ for Bitcoin Realized Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of $H(q)$ spectra.
@@
Table A.10: Summary of Log-Log Scaling Exponents for Bitcoin Realized Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of scaling coefficients.
@@
Table A.11: Summary of Wavelet Leaders Scaling Exponents for Bitcoin Realized Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) for each case.
@@
Table A.12: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, using the noise-robust realized volatility $\widetilde{RV}^{\Delta}_{t}$.